Fermi National Laboratory


Accelerator Update for March 23, 2005 - March 25, 2005

Wednesday March 23
The day shift began with Operations monitoring store 4056, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:20 AM, Operations reported that a Linac RF station (KRF7) had tripped off twice. An expert investigated and discovered contractors, who were pulling cables, were using radios to communicate and that the trips were caused by them keying their radios.

Operations terminated store 4056 at 9:03 AM.

At 9:07 AM, Operators reported more KRF7 problems. Experts investigated and said the station was having LCW regulation problems due to the recent trips.

Operations put the Tevatron (TeV) into shot setup at 9:23 AM.

At 9:44 AM, Operations reported that KRF7 had tripped off due to a small internal LCW leak. Experts had it repaired by 10:31 AM, which allowed shot setup to continue.

Operations established an Accumulator only store 4057 at 12:51 PM with an initial luminosity of 77.77E30.

MCenter resumed taking beam at 2:48 PM.

Operations put the Recycler into shot setup at 2:55 PM to stash antiprotons. Any shot setup halts beam to MCenter.

Operations reported at 3:06 PM that before the Recycler was put into shot setup, Main Injector (MI)/Pbar set a new stacking record with a one hour rate of 15.48mA.

Pbar resumed stacking at 4:06 PM.

An MI expert began a slip stacking study at 4:32 PM. The study lasted for about two hours.

MCenter resumed taking beam at 4:41 PM.

At 5:47 PM, a NuMI correction magnet power supply tripped off. Operators couldn't reset the supply. An expert came in and returned the supply to service about forty-five minutes later.

NuMI experts finished using the beam for the night at 9:11 PM. NuMI has a cooling system leak in its target as reported on Sunday night, 3/20/05. Experts are trying to determine how much beam NuMI can take with the present target while technicians prepare a second target.

Thursday March 24
The midnight shift began with Operations monitoring store 4057 and the Recycler stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 6:05 AM, LRF5 tripped off just as Operations received an alarm that the vacuum valves closed for KRF3-5. Operators investigated, found the vacuum good, and opened the valves. Operators turned on LRF5 locally and slowly brought up the current. The beam was off for about ten minutes.

At 8:21 AM, the Meson target train closed LCW system tripped off. An operator contacted the water system experts. The MCenter beam is now off. The expert estimated that the repairs would finish around 4 PM.

At 11:35 AM, antiprotons stopped making it to the Accumulator. Operators investigated and discovered that the injection shutter was closed due to a motor controller. A reset resolved the problem.

Operations terminated store 4057 at 12:07 PM and then put the TeV into shot setup.

Operations established store 4058 at 2:27 PM with an initial luminosity of 80.04E30. Pbar resumed stacking.

Technicians working on the Meson Target Train LCW system reported at 3:33 PM that they had completed their work.

Operations began stashing antiprotons into the Recycler at 6:45 PM. The stashing ended at 7:53 PM and Pbar began a stacking study. Operations halted beam to MiniBooNE for the study.

The Pbar stacking study ended at 10:01 PM. Pbar experts established a new stacking record with a rate of 16.2mA per hour. Operations resumed sending beam to MiniBooNE.

Operations established beam to MCenter at 11:26 PM.

Friday March 25
The midnight shift began with Operations monitoring store 4058 and the stash, with Pbar stacking, with Recycler experts conducting transfer studies, and with MiniBooNE and MCenter taking beam.

Recycler studies ended at 2 AM.

At 3:31 AM, BRF16 tripped off for the first of three times in three hours. It reset okay each time.

At 5:21 AM, the MI beam permit tripped off due to a power supply fault (KPS4A&B). The supplies reset okay.

The Plan for Wednesday
The plan for today is to terminate store 4058 around 1 PM and then put the TeV into shot setup. The plan for the rest of the weekend is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/25/2005   email Fermilab